OpenStudy (anonymous):

Please, help me solve x^2 f(x) = x+ integral_1^x (t f(t))/(t+1) dt, x>=1

OpenStudy (stamp):

not sure what that is supposed to be

OpenStudy (anonymous):

\[x^2f(x)=x+\int_1^x \frac{t\;f(t)}{t+1}dt,\;x\ge1\] I don't know either. Solving for f(x)? Finding f '(x)? @cke8, need more details.

OpenStudy (experimentx):

differentiate it and get differential equation and solve it.

OpenStudy (anonymous):

The only thing I know is that I should solve that integration equation but I don't know how! Do you?

OpenStudy (abb0t):

I believe that you're being tested the Fundamental Theorem of Calculus Pt. i

OpenStudy (anonymous):

Find \[f(x)\] that's about he only they can ask here.

OpenStudy (anonymous):

Here, first apply the derivative to everything with respect to x so, \[ 2xf(x)+2x^2f\prime(x)=1+\frac{xf(x)}{x+1}\] Next regroup so, \[ 2x^2f\prime(x)-1=\frac{xf(x)}{x+1}-2xf(x)\] \[ 2x^2f\prime(x)-1=\left( \frac{x}{x+1}-2x \right) f(x)\] \[ \frac{ 2x^2f\prime(x)-1}{\left( \frac{x}{x+1}-2x \right) }=f(x)\] And there you go, if they ask for the derivative jut solve for it.

OpenStudy (anonymous):

Thanks a lot!